You are working with an insurance company and need to take a survey of your clients to see if they would like the opportunity to buy health insurance for their pets. If you know you want a 95% confidence interval with a margin of error of 0.03, what how many clients do you need to survey?
Answer #1

The sample size formula is given as below:

n = p*q*(Z/E)^2

We are not given the estimate for population proportion, so we take

p = 0.5

q = 1 – p = 0.5

Confidence level = 95%

Critical Z value = 1.96

(by using z-table)

Margin of error = E = 0.03

The sample size is given as below:

n = p*q*(Z/E)^2

n = 0.5*0.5*(1.96/0.03)^2

n = 1067.111

n = 1068

Required sample size = 1068